Mez666:

Question #8 (there is only 1 cylinder)

Mez666:

1 attachment
AZ:

So you know how our formula is \( SA=2\pi rh+2\pi r^2\) It has \( 2\pi r^2\) because we have two circular bases your question says that the petri dish has no lid which means we only have one circle base so the surface area for the petri dish is going to be \( SA=2\pi rh+\pi r^2\)

Mez666:

12566.4

AZ:

That is correct!

Mez666:

Yay!